OpenStudy (anonymous):

Please confirm my answer? Solve for the unknown part of the triangle, a = 12, b = 15, c = 18 , if parts exist. C = 92 degrees 20'

OpenStudy (anonymous):

You want to know the angles?

OpenStudy (anonymous):

Well I need to know if I got the right answer (92 degrees 20") for C

OpenStudy (anonymous):

How did you get it?

OpenStudy (anonymous):

I used the law of cosines and I get that cosine of C = 0.25 http://en.wikipedia.org/wiki/Law_of_cosines And 92 doesn't seem plausible

OpenStudy (lgbasallote):

dont you have a figure to show?

OpenStudy (lgbasallote):

if i remember my trig right there isnt a formula for solving an angle given 3 sides is there?

OpenStudy (anonymous):

>>> a = 12 >>> b = 15 >>> c = 18 >>> (c**2 - b**2 - a**2)/(-2.0*a*b) 0.125 sorry 0.125 is the cosine of the angle @lgbasallote see above

OpenStudy (anonymous):

Well I got close to 90. Here are my other options 61°30' 65°20' 92°20' 82°50' And no I do not @Igba

OpenStudy (lgbasallote):

a 0.125 angle is very hard to believe though.....

OpenStudy (anonymous):

working on it, I think you're probably right after all hang on

OpenStudy (anonymous):

Thanks :)

OpenStudy (lgbasallote):

wiki says you should do \[\theta = \cos^{-1} \left(\frac{a^2 + b^2 - c^2}{2ab}\right)\]

OpenStudy (lgbasallote):

i think telliot forgot to arc cos or something

OpenStudy (lgbasallote):

and i think if you got an answer that's in the choices then it's most probably right

OpenStudy (anonymous):

cosine of 83 is about 0.122 @lgbasallote yes that is the law of cosines

OpenStudy (anonymous):

I'd go with 82°50'

OpenStudy (lgbasallote):

i got 82.82 from wolfram...

OpenStudy (anonymous):

Hmm, can you please give me the link to wolfram?

OpenStudy (anonymous):

Thanks!